OpenStudy (samsan9):

what are the coordinates of the vertex of the function below? Write in the form (x,y)? y-9=-6(x-1)^2

OpenStudy (kohai):

So to get this into a form where you can find the vertex, you need to get y by itself. This is simple enough, you add 9 to both sides. Your question now becomes: y=-6(x-1)^2+9 y=a(x-h)^2+k Where (h, k) is your vertex So your vertex is going to be (1, 9) Make sense?

OpenStudy (samsan9):

So whenever I get these problems i just use the coordinates from the formula in the function?

OpenStudy (kohai):

Yes, your vertex will always be (h, k). Notice that the vertex is (1, 9) here even though you're subtracting the 1 -- if it was (x+1)^2, the vertex would be (-1, 9).

OpenStudy (samsan9):

cool thank you

OpenStudy (kohai):

You're very welcome